Malcolm’s mother is 27 years older than he is and their ages are both two-digit numbers. If Malcolm swaps the digits in his age, he gets his mother’s age.
a How old is Malcolm if the sum of the digits in his age is 5?
b What is the relationship between the digits in Malcolm’s age if the sum of the digits is
unknown.
c If the sum of the digits in Malcolm’s two-digit age is unknown, how many possible ages
could he be? What are these ages?

Answer by mananth(16946)

Malcolm’s mother is 27 years older than he is and their ages are both two-digit numbers. If Malcolm swaps the digits in his age, he gets his mother’s age.
a How old is Malcolm if the sum of the digits in his age is 5?
b What is the relationship between the digits in Malcolm’s age if the sum of the digits is
unknown.
c If the sum of the digits in Malcolm’s two-digit age is unknown, how many possible ages
could he be? What are these ages?
let Malcolm's age be 10x+y
mothers age is 10x+y+27
equation after swapping the digits
10y+x = 10x+y+27
9y-9x=27
y-x=3
x+y=5
add the equations
2y=8
y=4
therefore x= 1
Malcom's age = 10x +y = 14
mother's age = 14+27 = 41